OpenStudy (anonymous):

h(x) = absolute value of (x-1) divided by (x+2). How do I write this function as a piecewise function? What are the horizontal asymptotes?

OpenStudy (gamer56):

What do you think it is first?

OpenStudy (anonymous):

\[h(x)=\frac{|x-1|}{x+2}\] like that?

OpenStudy (anonymous):

if that is it we can continue

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok we start with the definition of \(|x-1|\) which is \[ |x-1| = \left\{\begin{array}{rcc} -x+1 & \text{if} & x<1 \\ x-1& \text{if} & x\ge1 \end{array} \right.\]

OpenStudy (anonymous):

in more simple english it is itself if x - 1 is positive, and its opposite if x -1 is negative now all we need is to divide by \(x+2\)

OpenStudy (anonymous):

\[h(x) =\frac{|x-1|}{x+2}= \left\{\begin{array}{rcc} \frac{1-x}{x+2}& \text{if} & x<1 \\ \frac{x+1}{x+2}& \text{if} & x\geq 1 \end{array} \right.\]

OpenStudy (anonymous):

horizontal asymptotes next ?

OpenStudy (anonymous):

dont you mean to say (x-1)/(x+2) if x is greater than or equal to 1, and not (x+1) as the numerator?

OpenStudy (anonymous):

and yes, horizontal asymptotes i need help with. How would I go about solving?